Solve the following differential equations: (x^3+y^3)=(xy^2)dy/dx

1
Expert's answer
2022-05-11T12:48:59-0400

Solution

Given differential equation may be rewritten in the form

x3 + y3 = xy2dy/dx  =>  x2/y2 + y/x = dy/dx  

Let z=y/x  =>  y = z*x  =>  dy/dx = (dz/dx)x + z

From this and reduced equation

(dz/dx)x + z = 1/z2 + z  =>  dz/dx = 1/(xz2)  =>  z2dz = dx/x  =>

"\int{z^2dz}=\int\frac{dx}{x}" =>   "\frac{1}{3}z^3=ln\left|x\right|-C"   =>  (y/x)3 = 3ln|x| - 3C  =>  

"y(x)=x\sqrt[3]{3(ln\left|x\right|-C)}"


Answer

"y(x)=x\sqrt[3]{3(ln\left|x\right|-C)}"